OpenStudy (alyssakin):

(2^6)^x = 1 (5^0)^x = 1 Solve for x on both.

OpenStudy (anonymous):

\[(2^6)^x=2^{6x}\] and \(2^0=1\) set \[6x=0\] solve for \(x\)

OpenStudy (anonymous):

for the second one, \[5^0=1\] so it is true no matter what \(x\) is

OpenStudy (alyssakin):

OK, that makes sense now. Thank you.

OpenStudy (anonymous):

yw